OpenStudy (anonymous):

REALLY NEED HELP!!! PLEASE HELP ME!! For each equation, choose the statement that describes its solution.If applicable, give the solution. -3(y+1)+7y= 4(y+1)-5

OpenStudy (anonymous):

I'll Help

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sorry love to help but GTG :(

TheSmartOne (thesmartone):

First distribute the -3 inside the paranthesis and also the 4...

TheSmartOne (thesmartone):

You should get -3y-1+7y=4y+4-5

TheSmartOne (thesmartone):

4y-1=4y-1 0=0

TheSmartOne (thesmartone):

Oh wait I didn't distribue...

TheSmartOne (thesmartone):

-3y-3+7y=4y+4-5 4y-3=4y-1

TheSmartOne (thesmartone):

So yes it would be no solution..

TheSmartOne (thesmartone):

Becasue \[0\neq2\]

TheSmartOne (thesmartone):

Ok.. Let me see...

TheSmartOne (thesmartone):

I am not really sure... I am getting no solutions...

TheSmartOne (thesmartone):

Let me plug in numbers and see if it is all real numbers...

TheSmartOne (thesmartone):

Yes it is all real numbers...

TheSmartOne (thesmartone):

@babyblue123

TheSmartOne (thesmartone):

I am here... ^-^

OpenStudy (librarian):

Im here to sorry, Lately life has been really chaotic /=

OpenStudy (librarian):

I wil do my best (=

OpenStudy (librarian):

Here you go: http://www.algebra.com/services/rendering/simplifier.mpl

TheSmartOne (thesmartone):

It is all real solutions...

OpenStudy (librarian):

Go to that site and type in your problem and click the button and it will not only solve it but it will show you how to swell (=

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/ojwo8pbps9

TheSmartOne (thesmartone):

V=-2 https://www.desmos.com/calculator/vpq3oodgon

TheSmartOne (thesmartone):

@babyblue123

OpenStudy (anonymous):

For his exercise today, Eric plans to both run and swim. Let r be the number of laps he runs and let s be the number of laps he swims.Each lap he runs takes him 4 minutes, and each lap he swims takes him 3 minutes. He wants to exercise for more than 30 minutes today. Using the values and variables given, write an inequality describing this.

OpenStudy (anonymous):

@TheSmartOne help me with thiss

OpenStudy (anonymous):

pleaseee @TheSmartOne

TheSmartOne (thesmartone):

\[4r+3s \ge30\]